Math 291 Lab 9 Due Monday March 26, 2018

Instructions: Turn in your lab by emailing it to fagerstrom@mnstate.edu. You should email your raw TEX (.tex) file, your .bib file, your Maple (.mw) files, your .eps file, and your compiled document (in .ps format). Note: Make sure that you include at least these six files, along with any other picture files that you included with the includegraphics command.

You will be graded on both your raw TEX code and the accuracy of your compiled document. Name your file in the format of JaneDoeLab9.tex

Create a document that has at least:

(a) Bibliograph

(i) A bibliography that has at least four items (positioned at the end of your document, as usual for bibliographies).

(ii) The bibliography should start on a new page.

(iii) At least a paragraph or so of writing that refers to three of the items in your bibliography.

(iv) At least one item should appear in your bibliography that you have not cited in your document. [Use the \nocite command from page 13 of 27 of the lecture.]

(b) Includes five separate graphs:

(i) Includes an encapsulated postscript graph of a three-dimensional figure that was created in Maple.

(ii) Includes the graph of f(x) = 3x^{2} − 2x + 5 that was created in Maple and then downloaded as an .eps file. Make sure that you choose an appropriate domain/range in Maple.

(iii) Includes the graph of g(x) = 5x − 3 that was created using the psline command and the contents of the “smallgraph” file. Note: Make sure that you choose the starting and ending points of your line so that the line extends to the edges of the background grid and no further. Use arrowheads on both sides of the line, and make sure that they are big enough to be seen.

(iv) Includes the graph of some curve, of your choosing, using the pscurve command and the contents of the “largegraph” file. End your curve with an open circle on one end and a closed circle on the other end. Make sure that your curve is actually curved, and not a straight line.

(v) Includes the graphs of a circle, of an ellipse, and of a polygon with at least five sides, in a single pspicture, using the pscircle, psellipse, and pspolygon commands.